# A Simple Baseline for Supervised Surround-view Depth Estimation

### v3

@@ Abstract
Depth estimation has been widely studied and serves as the fundamental step of 3D perception for intelligent vehicles. Though significant progress has been made in monocular depth estimation in the past decades, these attempts are mainly conducted on the KITTI benchmark with only front-view cameras, which ignores the correlations across surround-view cameras. In this paper, we propose S3Depth, a Simple Baseline for Supervised Surround-view Depth Estimation, to jointly predict the depth maps across multiple surrounding cameras. Specifically, we employ a global-to-local feature extraction module that combines CNN with transformer layers for enriched representations. Further, the Adjacent-view Attention mechanism is proposed to enable the intra-view and inter-view feature propagation. The former is achieved by the self-attention module within each view, while the latter is realized by the adjacent attention module, which computes the attention across multi-cameras to exchange the multi-scale representations across surround-view feature maps. Extensive experiments show that our method achieves superior performance over existing state-of-the-art methods on both DDAD and nuScenes datasets.
@@ I Introduction
The rapid rise in Intelligent Vehicles (IVs) has captured global interest [1, 2, 3, 4], driven by their growing appeal for enhanced convenience, safety benefits, and commercial potential. With the rapid development of autonomous driving technology in recent years, depth estimation has received increasingly more attention. As a bridge connecting the 2D images and the 3D environments, depth estimation aims to recover the dense 3D representations from 2D images. Monocular depth estimation [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15] aims to get the depth map of a single image, which either uses LiDAR points as the target depth or learns from monocular videos in a self-supervised manner. Although significant progress has been made for monocular depth estimation in the past decade, existing methods are mainly performed on only a few driving scenes, mostly KITTI datasets [16]. Stereo matching [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32] attempts to predict the disparity between the left and right camera images, where a large overlap between the stereo images are essentially assumed. Though these methods achieve impressive performance for depth estimation, their receptive fields are limited to a small fraction of the entire 3D environment (usually less than $180^{\circ}$).
Although recently released datasets contain consistent multi-camera data that capture the full $360^{\circ}$ point cloud [33, 34], existing methods still handle the problems as monocular [35, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14] and ignore the multi-view interaction. Multi-View Stereo [36, 37, 38, 37, 39] is an effective method to predict a depth map by a series of calibrated images surrounding an object, but it needs large overlapping areas between cameras.
FSM [40] firstly extends self-supervised learning from monocular and stereo to the general multi-camera setting and introduces spatial-temporal contexts and pose consistency constraints, which boost performance in multi-camera settings. Following FSM [40], SurroundDepth [41] proposes to jointly process all surrounding views by the module of cross-view attention to get high-quality depth maps. In addition, SurroundDepth [41] is able to get the depths with real-world scales by the scale-aware SfM pretraining. However, the performance of these methods is insufficient for practical applications.
@@
Multi-view 3D object detection in Bird’s-Eye-View (BEV) has received great attention in recent years because it has a much lower computational cost for autonomous driving than LiDAR-based solutions. Estimating accurate depth is essential for multi-view 3D object detection. However, it remains challenging to predict depth in surrounding images without overlapping areas. BEVDepth [42] provides a quantitative and qualitative analysis of depth estimation within 3D detectors, and boosts the performance of 3D detectors by using ground-truth depth. This reveals that accurate depth is helpful for multi-view 3D object detection. Based on BEVDepth [42] and inspired by the success of the MVS and monocular depth estimation methods [36, 43, 44, 45, 46], STS [47] proposes the temporal stereo technique called Surround-view Temporal Stereo (STS), which leverages the geometric correspondence across frames over time. Different from BEVDepth [42] which predicts depth from a single image, BEVStereo [48] further predicts both the monocular and stereo depths to get better BEV representations for 3D object detection. Although these methods have explored the effect of depth estimation on BEV 3D object detection, there is no baseline for supervised surround-view depth estimation.
In this paper, we propose a Simple Baseline for Supervised Surround-view Depth estimation, called S3Depth. Specifically, we first employ a shared encoder that can extract global-to-local features by combining CNN with transformer layers. To learn adjacent-view interactions in surrounding views, we adopt the alternating attention mechanism: self-attention computes attention in the same view, while adjacent attention computes attention between adjacent views. Finally, different from monocular depth estimation, a Depth Head is used to predict all surround-view depth maps at the same time. We conduct extensive experiments on the challenging datasets DDAD [33] and nuScenes [34] for multi-camera depth estimation. Experimental results show that the adjacent-view feature interaction boosts the performance of depth estimation.
@@
  • For the first time, we propose a simple baseline for supervised surround-view depth estimation, to jointly predict depth maps in a wide-baseline $360\degree$ multi-camera setting.
  • We use the joint convolutional attention and transformer block for surround-view depth estimation to model local and global contexts.
  • We propose Adjacent-view Attention mechanism which can be used to enable the intra-view and inter-view feature propagation in this novel setting.
  • Our S3Depth yields substantial improvements to depth estimation when compared to state-of-the-art methods on two publicly available multi-camera datasets: DDAD [33] and nuScenes [34].
@@ II Related Work
Monocular Depth Estimation. Monocular depth estimation has been studied in computer vision for a long time and plays a crucial role in 3D reconstruction, autonomous driving cars, and robot navigation. Supervised approaches take a single image and use datasets collected by laser scanners [16] or other range sensors [49] as target depth labels in training, as shown in Figure 1(a). Although learning-based depth estimation methods [5, 6, 7, 8, 9, 10, 11, 12, 13, 50, 35, 51, 52, 53, 54] achieve impressive results when compared with traditional machine learning baselines [7, 55, 56], these methods need additional sensors which are used to collect “ground-truth” data. Besides, these methods are designed for the KITTI datasets [16] whose “ground-truth” depth maps are dense. Moreover, the “ground-truth” data may be sparsity and not be very reliable due to high noise during data collection in practical applications. Thus, self-supervised methods [57, 58, 25, 59, 60, 61, 62, 63, 64] explore the routes of jointly learning monocular depth and ego-motion from unlabelled sequential data, as shown in Figure 1(b). Zhou et al [57] convert the monocular depth estimation into a task of view synthesis, using the photometric loss as the supervisory signal to train networks to separately predict poses and depths. To enhance the quality of predictions, Monodepth2  [60] presents an auto-masking loss and a minimum reprojection loss, which can handle occlusions and moving objects. Recently, inspired by the success of transformer [65, 66, 67, 68, 69, 45], MonoVit  [14] further improves depth estimation accuracy by combining plain convolutions with transformer blocks. However, these approaches are designed for monocular depth estimation (typically less than $180^{\circ}$), consistent surround-view depth estimation would extend these approaches to the full $360^{\circ}$ space.
Figure 2: The architecture of our proposed S3Depth. Our S3Depth consists of three part: Feature Extraction, Adjacent-view Attention and Depth Head. For feature extraction, both transformer and convolutional layers are adopted to enhance the feature modeling and depth inferring. For Adjacent-view Attention, the alternating self-attention and adjacent-attention are adopted for $Z$ times to fully exchange the information between current view and adjacent view. Finally, Depth Head predicts the all-view depth map at the same time.
@@
Stereo Matching. Stereo matching attempts to predict disparity between the left and right camera image, which can be used to estimate depth, as shown in Figure 1(c). In general, the stereo depth estimation task treats the 2D matching problem as a 1D disparity search. Traditional methods [70, 71] utilize hand-crafted features to computer matching cost, which tends to fail on those textureless and repetitive regions in the images. In recent years, with the support of large synthetic datasets [17, 72, 73], the end-to-end network has achieved impressive performance on stereo matching. One line of networks [17, 18, 19, 20, 21, 22, 23] match features by dot-product correlation, which only uses 2D convolutions. Another line of networks [24, 25, 26, 27, 28, 29] build 4D features volume and learn to aggregate the matching cost by 3D convolutions. Besides, Li et al [30] take advantage of the transformer architecture, which has advantages in feature matching and relaxes the limitation of a fixed disparity range, and presents a new end-to-end-trained network called STTR. However, a large overlap between left and right images is needed to train these models. Our proposed method is intended for consistent multi-camera configurations with minimal image overlap.
@@
Surround-view Depth Estimation. Based on the way of supervision, surround-view depth estimation methods can be divided into self-supervised and supervised. To our knowledge, FSM [40] firstly extends self-supervised monocular depth estimation to a large-baseline multi-camera setting. Following FSM [40], SurroundDepth [41] recovers the real-world scales by pretraining the depth network with sparse depths obtained from SFM [74] and learn to capture the surround-view interactions by the module of cross-view attention, as shown in Figure 1(e). Cross-view transformer employed in SurroundDepth [41] combines six views through stitching and is implemented using self-attention layers. For the task of supervised multi-view depth estimation, it is an effective method to build cost volume [36, 37, 38, 37, 39], namely, multi-view stereo (MVS). MVS aims to estimate the depth map by a series of calibrated images surrounding an object, and it is a one-to-many feature matching task, as shown in Figure 1(d). Different from MVS, our model is designed to estimate all depth maps of surrounding multi-view images with small overlapping areas between cameras.
Our multi-camera setting is related to the works [42, 48, 47], where consistent multi-camera depth maps are used to predict 3D bounding boxes. BEVDepth [42] introduces a depth correction module supervised explicitly by LiDAR, which encodes the camera parameters and is helpful for 3D object detection. Inspired by the success of the MVS and monocular depth estimation methods [36, 43, 44, 45, 46], STS [47] proposes the temporal stereo technique called Surround-view Temporal Stereo (STS), which is used to facilitate 3D detection. To get better BEV representations for 3D object detection, BEVStereo [48] enhances depth estimation with dynamic temporal stereo, which is able to handle complex outdoor scenarios. Although those methods [42, 47, 48] have explored the effect of depth estimation on 3D object detection, there has been a lack of a baseline for supervised surround-view depth estimation. Therefore, it is necessary to establish a simple yet effective baseline for supervised surround-view depth estimation in order to evaluate the performance of more advanced methods and facilitate the development of this research field.
@@ III-A Problem Formulation
Given ${N}$ views input samples ${I}=\{{I}^{1},{I}^{2},\cdots{I}^{N}\}$, the surround-view depth model $f(\cdot)$ is expected to predict the depth map ${D}^{pred}=\{{D}^{pred}_{1},{D}^{pred}_{2},\cdots{D}^{pred}_{N}\}$, where ${I}^{i}\in\mathbb{R}^{H\times W\times 3}$ denotes the surround-view RGB images, ${D}^{pred}_{i}\in\mathbb{R}^{H\times W}$ denotes the predicted depth map and ${H}$, ${W}$ stand for the height and width of input images, respectively. In S3Depth, the only supervised signal comes from LiDAR. Different from monocular depth estimation, there are crucial overlaps across the surround-view images, which are very useful and can be imported for the model to understand the scene as a whole. Based on that, we introduce a simple baseline S3Depth to predict $360^{\circ}$ view depth at the same time. By taking advantage of the adjacent-view information interactions, S3Depth can achieve better performance of every view than view-dependent estimations.
@@ III-B Overview
Figure 2 shows our proposed S3Depth architecture. Our purpose is to build a powerful baseline for dense depth prediction of surround-view images, so we employ a multi-stage model [67] to get enriched representations. Specifically, the baseline consists of three parts: Feature Extraction, Adjacent-view Attention, and Depth Head. (1) Feature extraction is a shared backbone that extracts 2D features from $N$ views of input images. (2) The adjacent-view Attention architecture is leveraged to exchange adjacent-view information of features. Specifically, the alternating attention mechanism is adopted: self-attention computes attention to features in the same view, while adjacent-attention computes attention to features in the adjacent views to exchange the adjacent multi-scale representations at every scale. (3) Different from monocular depth estimation, the Depth Head predicts the surround-view depth map at the same time.
Figure 3: The module combining CNN and Transformer layers. Trans stands for Transformer.
@@ III-C Feature Extractor
Image Encoder. Recently, the attention mechanism has already proven to be effective in various computer vision tasks, such as image classification [66], object detection [75] and depth estimation [68, 69, 45, 14]. Benefiting from the inductive bias, the transformer allows the model to focus on the important parts of the feature. Although the transformer can model global context perception, it may ignore local relationships. On the contrary, CNN can take advantage of local connectivity. Therefore, we adopt a module that combines CNN with transformer [67] within S3Depth, as shown in Figure 3, which can learn to interact between local and global features for enriched representations in five stages. Specifically, in the first stage, a Conv-stem block is adopted to extract 2D features $F_{1}\in\mathbb{R}^{N\times C_{F}\times\frac{H}{2}\times\frac{W}{2}}$, where $H$ and $W$ indicate the height and width of images, $N$ represents views of input images and $C_{F}$ stands for the feature channel number. For the next four stages $i\in\{2,3,4,5\}$, the joint convolution attention and transformer blocks are used to extract a features pyramid $\{F_{2}\in\mathbb{R}^{N\times C_{F}\times\frac{H}{2}\times\frac{W}{2}},F_{3}\in\mathbb{R}^{N\times C_{F}\times\frac{H}{4}\times\frac{W}{4}},F_{4}\in\mathbb{R}^{N\times C_{F}\times\frac{H}{8}\times\frac{W}{8}},F_{5}\in\mathbb{R}^{N\times C_{F}\times\frac{H}{16}\times\frac{W}{16}}\}$, which can represent multi-scale features.
@@ III-D Adjacent-view Attention
The detailed structure of the adjacent-view attention architecture is provided in Figure 2. Different from SurroundDepth [41], which utilizes a cross-view transformer that combines six views through stitching and implements it using self-attention layers, we adopt the alternating self-attention and adjacent-attention. Self-attention extracts features within the current view and adjacent-attention is used to exchange representations across adjacent view features. Let $F_{ij}\in\mathbb{R}^{N\times C_{F}\times\frac{H}{n}\times\frac{W}{n}},i=1,2,3,4,5,j=1,2,3,4,5,6$ be the feature maps obtained from the $j$-th view of the $i$-th stage, where $H$ and $W$ indicates the height and width of images, $N$ represents views of input images, $C_{F}$ stands for the feature channel number. For every stage, multi-head attention is used to increase the expressiveness of the features. The attention mechanism uses the dot product similarity to calculate the attention between the $query$ vectors and the $key$ vectors, and then weigh the $value$ vectors. For self-attention, the $\mathcal{Q},\mathcal{K},\mathcal{V}$ are computed from the same view using feature descriptors $F_{ij}$ as input, where $W_{\mathcal{Q}},W_{\mathcal{K}},W_{\mathcal{V}}\in\mathbb{R}^{C\times C}$, $b_{\mathcal{Q}},b_{\mathcal{K}},b_{\mathcal{V}}\in\mathbb{R}^{C}$ and C represents the embedding dimension respectively:
$\begin{split}Q_{i,j}&=W_{\mathcal{Q}}F_{i,j}+b_{\mathcal{Q}}\\ K_{i,j}&=W_{\mathcal{K}}F_{i,j}+b_{\mathcal{K}}\\ V_{i,j}&=W_{\mathcal{V}}F_{i,j}+b_{\mathcal{V}}\\ F^{sf}_{i,j}&=\text{softmax}(\frac{Q_{i,j}^{T}K_{i,j}}{\sqrt{C}}){V_{i,j}}\end{split}$ (1)
For adjacent-attention, $\mathcal{K},\mathcal{V}$ are computed from the current-view feature, while $\mathcal{Q}$ is computed from the adjacent-view feature, where $x$ refers to adjacent views ($x\in{j-1,j+1}$):
$\begin{split}Q_{i,x}&=W_{\mathcal{Q}}F^{sf}_{i,x}+b_{\mathcal{Q}}\\ K_{i,j}&=W_{\mathcal{K}}F^{sf}_{i,j}+b_{\mathcal{K}}\\ V_{i,j}&=W_{\mathcal{V}}F^{sf}_{i,j}+b_{\mathcal{V}}\\ F^{adj}_{i,j}&=\text{softmax}(\frac{Q_{i,x}^{T}K_{i,j}}{\sqrt{C}}){V_{i,j}}\end{split}$ (2)
@@
To achieve a comprehensive understanding and powerful representation features at the same location at different scales are used together to predict the final depth. By taking advantage of the adjacent-view attention, features can interact with adjacent views to compensate for erroneous estimations. Let $F^{out}$ denote the features from the adjacent-attention, the final output can be represented as follows:
$\begin{split}F^{final}&=\sigma(F^{out}){D_{max}}\end{split}$ (3)
where $\sigma$ stands for the $sigmoid$ activation function and $D_{max}$ denotes the max distance.

### v4

# Adjacent-view Transformers for Supervised Surround-view Depth Estimation

@@ Abstract
Depth estimation has been widely studied and serves as the fundamental step of 3D perception for robotics and autonomous driving. Though significant progress has been made in monocular depth estimation in the past decades, these attempts are mainly conducted on the KITTI benchmark with only front-view cameras, which ignores the correlations across surround-view cameras. In this paper, we propose an Adjacent-View Transformer for Supervised Surround-view Depth estimation (AVT-SSDepth), to jointly predict the depth maps across multiple surrounding cameras. Specifically, we employ a global-to-local feature extraction module that combines CNN with transformer layers for enriched representations. Further, the adjacent-view attention mechanism is proposed to enable the intra-view and inter-view feature propagation. The former is achieved by the self-attention module within each view, while the latter is realized by the adjacent attention module, which computes the attention across multi-cameras to exchange the multi-scale representations across surround-view feature maps. In addition, AVT-SSDepth has strong cross-dataset generaliza- tion. Extensive experiments show that our method achieves superior performance over existing state-of-the-art methods on both DDAD and nuScenes datasets. Code is available at https://github.com/XiandaGuo/SSDepth.
11footnotetext: These authors contributed to the work equally.22footnotetext: Project Leader.
, 33footnotetext: Corresponding authors.
@@ I Introduction
With the rapid development of robotics and autonomous driving technology in recent years, depth estimation has received increasingly more attention. As a bridge connecting the 2D images and the 3D environments, depth estimation aims to recover the dense 3D representations from 2D images. Monocular depth estimation [1, 2, 3, 4] aims to get the depth map of a single image, which either uses LiDAR points as the target depth or learns from monocular videos in a self-supervised manner. Although significant progress has been made for monocular depth estimation in the past decade, existing methods are mainly performed on only a few driving scenes, mostly KITTI datasets [5]. Stereo matching [6, 7, 8] attempts to predict the disparity between the left and right camera images, where a large overlap between the stereo images is assumed. Though these methods achieve impressive performance for depth estimation, their receptive fields are limited to a small fraction of the entire 3D environment (usually less than $180^{\circ}$).
Although recently released datasets contain consistent multi-camera data that capture the full $360^{\circ}$ point cloud [9, 10], existing methods still handle the problems as monocular [11, 12, 13, 14, 15] and ignore the multi-view interaction. Multi-View Stereo [16, 17, 18, 17, 19] is an effective method to predict a depth map by a series of calibrated images surrounding an object, but it needs large overlapping areas between cameras. FSM [20] firstly extends self-supervised learning from monocular and stereo to the general multi-camera setting and introduces spatial-temporal contexts and pose consistency constraints, which boost performance in multi-camera settings. Following FSM [20], SurroundDepth [21] proposes to jointly process all surrounding views by the module of cross-view attention to get high-quality depth maps. In addition, SurroundDepth [21] can get the depths with real-world scales by the scale-aware SfM pretraining. However, the performance of these methods is insufficient for practical applications.
Multi-view 3D object detection in Bird’s-Eye-View (BEV) has received great attention recently because it has a much lower computational cost for autonomous driving than LiDAR-based solutions. Estimating accurate depth is essential for multi-view 3D object detection. However, it remains challenging to predict depth in surrounding images without overlapping areas. BEVDepth [22] provides a quantitative and qualitative analysis of depth estimation within 3D detectors and boosts the performance of 3D detectors by using ground-truth depth. Although these methods have explored the effect of depth estimation on BEV 3D object detection, there is no baseline for supervised surround-view depth estimation. Therefore, it is time to build a baseline for supervised surround-view depth estimation.
@@
In this paper, we propose Adjacent-View Transformers Supervised Surround-view Depth estimation, called AVT-SSDepth. Specifically, we first employ a shared encoder that can extract global-to-local features by combining CNN with transformer layers. To learn adjacent-view interactions in surrounding views, we adopt the alternating attention mechanism: self-attention computes attention in the same view, while adjacent attention computes attention between adjacent views. Finally, different from monocular depth estimation, a Depth Head is used to predict all surround-view depth maps at the same time. We conduct extensive experiments on the challenging datasets DDAD [9] and nuScenes [10] for multi-camera depth estimation.
@@
  • For the first time, we propose adjacent-view transformers for supervised surround-view depth estimation to jointly predict depth maps in a wide-baseline $360^{\circ}$ multi-camera setting.
  • We use the joint convolutional attention and transformer block for surround-view depth estimation to model local and global contexts. Furthermore, we propose adjacent-view attention mechanism, which can be used to enable the intra-view and inter-view feature propagation in this novel setting.
  • Our AVT-SSDepth yields substantial improvements to depth estimation when compared to state-of-the-art methods on DDAD [9] and nuScenes [10].
@@ II Related Work
Monocular Depth Estimation. Monocular depth estimation has been studied in computer vision for a long time and plays a crucial role in 3D reconstruction, autonomous driving cars, and robot navigation. Supervised approaches take a single image and use datasets collected by laser scanners [5] or other range sensors [23] as target depth labels in training, as shown in Fig.1(a). Although learning-based depth estimation methods [12, 1, 24, 25, 11, 26, 27, 28, 29] achieve impressive results when compared with traditional machine learning baselines [30, 31, 32], these methods need additional sensors which are used to collect “ground-truth" data. Besides, these methods are designed for the KITTI datasets [5] whose “ground-truth" depth maps are dense. Moreover, the “ground-truth" data may be sparse and not be very reliable due to high noise during data collection in practical applications. Thus, self-supervised methods [33, 34, 6, 35, 2, 36, 37, 38, 39] explore the routes of jointly learning monocular depth and ego-motion from unlabeled sequential data, as shown in Fig.1(b). Inspired by the success of transformer [40, 41, 42, 43, 44, 45], MonoVit [3] further improves depth estimation accuracy by combining plain convolutions with transformer blocks. Recently, some works have aimed to build a powerful foundation model, such as MiDaS [46], DepthAnything [28, 29], and Marigold [47]. However, these methods are designed for monocular depth estimation (typically less than $180^{\circ}$), consistent surround-view depth estimation would extend these approaches to the full $360^{\circ}$ space.
@@
Stereo Matching. Stereo matching predicts disparity between the left and right image, as shown in Fig.1(c). In general, the stereo depth estimation task treats the 2D matching problem as a 1D disparity search. Traditional methods [48, 49] utilize hand-crafted features to computer matching cost, which tends to fail on those textureless and repetitive regions in the images. In recent years, with the support of large synthetic datasets [50, 51, 52], the end-to-end network has achieved impressive performance on stereo matching. One line of networks [50, 53, 54, 55, 56, 57, 58] match features by dot-product correlation, which only uses 2D convolutions. Another line of networks [59, 6, 60, 61, 62, 63] builds 4D feature volume and learns to aggregate the matching cost by 3D convolutions. Besides, Li et al [64] take advantage of the transformer architecture, which has advantages in feature matching and relaxes the limitation of a fixed disparity range, called STTR. However, a large overlap between the left and right images is needed for stereo matching. Our proposed method is intended for consistent multi-camera configurations with minimal image overlap.
Figure 2: The architecture of our proposed AVT-SSDepth. Our AVT-SSDepth consists of three parts: Feature Extraction, Adjacent-view Attention, and Depth Head. For feature extraction, both transformer and convolutional layers are adopted to enhance the feature modeling and depth inferring. For Adjacent-view Attention, the alternating self-attention and adjacent-attention are adopted for $Z$ times to fully exchange the information between the current view and the adjacent view. Finally, Depth Head predicts the all-view depth map at the same time.
@@
Surround-view Depth Estimation. Based on the way of supervision, surround-view depth estimation methods can be divided into self-supervised and supervised. To our knowledge, FSM [20] first extends self-supervised monocular depth estimation to a large-baseline multi-camera setting. Following FSM [20], SurroundDepth [21] recovers the real-world scales by pretraining the depth network with sparse depths obtained from SFM [65] and learns to capture the surround-view interactions using self-attention layers, as shown in Fig.1(e). Recently, R3D3 [66] leverages both spatial and temporal cues for dense 3D reconstruction from multiple cameras in a self-supervised manner. For the task of supervised multi-view depth estimation, it is an effective method to build cost volume [16, 18, 17, 19, 67], namely, multi-view stereo (MVS). MVS aims to estimate the depth map by a series of calibrated images surrounding an object, and it is a one-to-many feature matching task, as shown in Fig.1(d). Different from MVS, our model is designed to estimate all depth maps of surrounding multi-view images with small overlapping areas between cameras.
@@ III-A Problem Formulation
Given N{N} views input samples ${I}=\{{I}^{1},{I}^{2},\cdots{I}^{N}\}$, the surround-view depth model $f(\cdot)$ is expected to predict the depth map ${D}^{pred}=\{{D}^{pred}_{1},{D}^{pred}_{2},\cdots{D}^{pred}_{N}\}$, where ${I}^{i}\in\mathbb{R}^{H\times W\times 3}$ denotes the surround-view RGB images, ${D}^{pred}_{i}\in\mathbb{R}^{H\times W}$ denotes the predicted depth map and H{H}, W{W} stand for the height and width of input images, respectively. In AVT-SSDepth, the only supervised signal comes from LiDAR. Different from monocular depth estimation, there are crucial overlaps across the surround-view images, which are very useful and can be imported for the model to understand the scene as a whole. Based on that, we introduce a simple baseline AVT-SSDepth to predict $360^{\circ}$ view depth at the same time. By taking advantage of the adjacent-view information interactions, AVT-SSDepth can achieve better performance of every view than view-dependent estimations.
@@ III-B Overview
Fig.2 shows our proposed AVT-SSDepth architecture. Our purpose is to build a powerful baseline for dense depth prediction of surround-view images, so we employ a multi-stage model [42] to get enriched representations. Specifically, the baseline consists of three parts: Feature Extraction, Adjacent-view Attention, and Depth Head. (1) Feature extraction is a shared backbone that extracts 2D features from $N$ views of input images. (2) The adjacent-view Attention architecture is leveraged to exchange adjacent-view information of features. Specifically, the alternating attention mechanism is adopted: self-attention computes attention to features in the same view, while adjacent attention computes attention to features in the adjacent views to exchange the adjacent multi-scale representations at every scale. (3) Different from monocular depth estimation, the Depth Head predicts the surround-view depth map at the same time.
@@ III-C Feature Extractor
Image Encoder. Recently, the attention mechanism has already proven to be effective in various computer vision tasks, such as image classification [41], object detection [68] and depth estimation [43, 44, 45, 3, 69]. Benefiting from the inductive bias, the transformer allows the model to focus on the important parts of the feature. Although the transformer can model global context perception, it may ignore local relationships. On the contrary, CNN can take advantage of local connectivity. Therefore, we adopt a module that combines CNN with transformer [42] within AVT-SSDepth, which can learn to interact between local and global features for enriched representations in five stages. Multiple surround-view images are concatenated and then fed into a feature extractor. Specifically, in the first stage, a Conv-stem block is adopted to extract 2D features $F_{1}\in\mathbb{R}^{N\times C_{F}\times\frac{H}{2}\times\frac{W}{2}}$, where $H$ and $W$ indicate the height and width of images, $N$ represents views of input images and $C_{F}$ stands for the feature channel number. $F_{1}$ denotes features of multiple views. For the next four stages $i\in\{2,3,4,5\}$, the joint convolution attention and transformer blocks are used to extract a features pyramid $\{F_{2}\in\mathbb{R}^{N\times C_{F}\times\frac{H}{2}\times\frac{W}{2}},F_{3}\in\mathbb{R}^{N\times C_{F}\times\frac{H}{4}\times\frac{W}{4}},F_{4}\in\mathbb{R}^{N\times C_{F}\times\frac{H}{8}\times\frac{W}{8}},F_{5}\in\mathbb{R}^{N\times C_{F}\times\frac{H}{16}\times\frac{W}{16}}\}$, which can represent multi-scale features.
@@ III-D Adjacent-view Attention
The detailed structure of the adjacent-view attention architecture is provided in Fig.2. Different from SurroundDepth [21], which utilizes a cross-view transformer that combines six views through stitching and implements it using self-attention layers, we adopt the alternating self-attention and adjacent-attention. Self-attention extracts features within the current view and adjacent-attention is used to exchange representations across adjacent view features. Let $F_{ij}\in\mathbb{R}^{N\times C_{F}\times\frac{H}{n}\times\frac{W}{n}},i=1,2,3,4,5,j=1,2,3,4,5,6$ be the feature maps obtained from the $j$-th view of the $i$-th stage, where $H$ and $W$ indicates the height and width of images, $N$ represents views of input images, $C_{F}$ stands for the feature channel number. For every stage, multi-head attention is used to increase the expressiveness of the features. The attention mechanism uses the dot product similarity to calculate the attention between the queryquery vectors and the keykey vectors, and then weigh the valuevalue vectors. For self-attention, the $\mathcal{Q},\mathcal{K},\mathcal{V}$ are computed from the same view using feature descriptors $F_{ij}$ as input, where $W_{\mathcal{Q}},W_{\mathcal{K}},W_{\mathcal{V}}\in\mathbb{R}^{C\times C}$, $b_{\mathcal{Q}},b_{\mathcal{K}},b_{\mathcal{V}}\in\mathbb{R}^{C}$ and C represents the embedding dimension respectively:
$Q_{i,j}=W_{\mathcal{Q}}F_{i,j}+b_{\mathcal{Q}},K_{i,j}=W_{\mathcal{K}}F_{i,j}+b_{\mathcal{K}},V_{i,j}=W_{\mathcal{V}}F_{i,j}+b_{\mathcal{V}},$ (1)
$\begin{split}F^{sf}_{i,j}&=\text{softmax}(\frac{Q_{i,j}^{T}K_{i,j}}{\sqrt{C}}){V_{i,j}}.\end{split}$ (2)
For adjacent-attention, $\mathcal{K},\mathcal{V}$ are computed from the current-view feature, while $\mathcal{Q}$ is computed from the adjacent-view feature, where $x$ refers to adjacent views ($x\in{j-1,j+1}$). Specifically, when $j$ equals 1, $x$ equals 6 and 2; Similarly, when $j$ equals 6, $x$ equals 5 and 1:
$Q_{i,x}=W_{\mathcal{Q}}F^{sf}_{i,x}+b_{\mathcal{Q}},$ (3)
$K_{i,j}=W_{\mathcal{K}}F^{sf}_{i,j}+b_{\mathcal{K}},$ (4)
$V_{i,j}=W_{\mathcal{V}}F^{sf}_{i,j}+b_{\mathcal{V}},$ (5)
$\begin{split}F^{adj}_{i,j}&=\text{softmax}(\frac{Q_{i,x}^{T}K_{i,j}}{\sqrt{C}}){V_{i,j}}.\end{split}$ (6)
@@
To achieve a comprehensive understanding and powerful representation features at the same location at different scales are used together to predict the final depth. By taking advantage of the adjacent-view attention, features can interact with adjacent views to compensate for erroneous estimations. The final output can be represented as follows:
$\begin{split}D^{pred}_{i}&=\sigma(F^{out}_{i}){D_{max}},\end{split}$ (7)
where $\sigma$ stands for the sigmoidsigmoid activation function and $D_{max}$ denotes the max distance. $F^{out}_{i}$ represents the upsampled features of the adjacent attention from each stage, bringing them back to the original image resolution.